1. Machinery was purchased on January 1 for $73,500.00. The machinery has an estimated life of seven years and an estimated salvage value of $9,000. Double-declining-balance depreciation for the second year would be (round calculations to the nearest dollar):

a.$16,000

b.$15,000

c.$14,500

d.$14,000

2. A machine with a cost of $52,300 has an estimated residual value of $3,262 and an estimated life of 5 years or 17,412 hours. What is the amount of depreciation for the second full year, using the double-declining-balance method?

a.$12,552

b.$19,615

c.$20,920

d.$10,460

3. A building with an appraisal value of $135,757 is made available at an offer price of $158,563. The purchaser acquires the property for $34,867 in cash, a 90-day note payable for $23,496, and a mortgage amounting to $58,384. The cost basis recorded in the buyer's accounting records to recognize this purchase is

a.$158,563

b.$135,757

c.$123,696

d.$116,747

4. A machine with a cost of $135,000 has an estimated residual value of $15,000 and an estimated life of 6 years or 15,000 hours. It is to be depreciated by the units-of-activity method. What is the amount of depreciation for the second full year, during which the machine was used 5,100 hours?

a.$51,000

b.$40,800

c.$25,800

d.$5,100

5. On December 31, Strike Company has decided to discard one of its batting cages. The equipment had an initial cost of $205,000 and has accumulated depreciation of $184,500. Depreciation has been recorded up to the end of the year. Which of the following will be included in the entry to record the disposal?

a.Equipment, credit, $205,000

b.Gain on Disposal of Asset, credit, $20,500

c.Loss on Disposal of Asset, debit, $184,500

d.Accumulated Depreciation, debit, $205,000

6. Equipment with a cost of $154,821 has an estimated residual value of $8,049 and an estimated life of 4 years or 15,860 hours. It is to be depreciated by the straight-line method. What is the amount of depreciation for the first full year, during which the equipment was used for 3,136 hours?

a.$36,693.00

b.$29,021.25

c.$38,705.25

d.$29,068.05

7.

Computer equipment was acquired at the beginning of the year at a cost of $58,800. It had an estimated residual value of $4,800 and an estimated useful life of five years. Determine the (a) depreciable cost, (b) straight-line rate, and (c) annual straight-line depreciation.

a. Depreciable cost $
b. Straight-line rate %
c. Annual straight-line depreciation $
Answer #1
1)
Depreciation rate under straight line = 1/7 years =14.285%
Depreciation under double decling method =14.285*2 =28.57%
Depreciation Schedule
Year opening balance Depreciation @28.57% Accumulated depreciation Closing balance
1 $                    73,500 $                               20,999 $                                       20,999 $         52,501
2 $                    52,501 $                               15,000 $                                       35,998 $         37,502
3 $                    37,502 $                               10,714 $                                       46,713 $         26,787
4 $                    26,787 $                                 7,653 $                                       54,366 $         19,134
5 $                    19,134 $                                 5,467 $                                       59,832 $         13,668
6 $                    13,668 $                                 3,905 $                                       63,737 $           9,763
7 $                       9,763 $                                 2,789 $                                       66,526 $           6,974
Correct Option: b
2) Depreciation rate under straight line = 1/5years =20%
Depreciation under double decling method =20%*2 =40%
Depreciation Schedule
Year opening balance Depreciation @40% Accumulated depreciation Closing balance
1 $                    52,300 $                               20,920 $                                       20,920 $         31,380
2 $                    31,380 $                               12,552 $                                       33,472 $         18,828
3 $                    18,828 $                                 7,531 $                                       41,003 $         11,297
4 $                    11,297 $                                 4,519 $                                       45,522 $           6,778
5 $                       6,778 $                                 2,711 $                                       48,233 $           4,067
Correct Option: a